Let f: I → R, where I is an open interval containing the point c, and let k ∈ R. Prove the following.


(a) f is differentiable at c with f ′(c) = k iff limh→0 [ f (c + h) – f (c)]/h = k.


*(b) If f is differentiable at c with f ′(c) = k, then limh→ 0 [ f (c + h) – f (c – h)]/2h = k.


(c) If f is differentiable at c with f ′(c) = k, then lim n →∞ n[f (c + 1/n) – f (c)] = k.


(d) Find counterexamples to show that the converses of parts (b) and (c) are not true.




The book is Steven R. Lay, Analysis with an introduction to proof.


1
Expert's answer
2022-02-21T17:45:16-0500

a) By the definition, "f" is differentiable at "c" with "f '(c)=k" iff

"f(x)=f(c)+k(x-c)+o(|x-c|),\,x\to c". Then, by denoting "x-c" as h, we have:

"f" is differentiable at "c" with "f '(c)=k" iff

"f(c+h)-f(c)=kh+o(|h|),\,h\to 0" iff

"\frac{f(c+h)-f(c)}{h}=k+\frac{o(|h|)}{h}=k+o(1)" iff

"\lim\limits_{h\to 0}\frac{f(c+h)-f(c)}{h}=k".


b) Using the result from a), we have

"\lim\limits_{h\to 0}\frac{f(c+h)-f(c)}{h}=k" and "\lim\limits_{h\to 0}\frac{f(c)-f(c-h)}{h}=\lim\limits_{h\to 0}\frac{f(c-h)-f(c)}{-h}=\lim\limits_{h\to 0}\frac{f(c+h)-f(c)}{h}=k".

Therefore,

"\lim\limits_{h\to 0}\frac{f(c+h)-f(c-h)}{2h}=\frac{1}{2}\lim\limits_{h\to 0}(\frac{f(c+h)-f(c)}{h}+\frac{f(c)-f(c-h)}{h})=\frac{1}{2}(k+k)=k".


c) If "f" is differentiable at "c" with "f'(c)=k" , then "\lim\limits_{h\to 0}\frac{f(c+h)-f(c)}{h}=k". That is, for all "\varepsilon>0" there exists "\delta>0" such that for all "h\in(-\delta,\delta)" we have "|\frac{f(c+h)-f(c)}{h}-k|<\varepsilon".

Put "N=[1/\delta]+1>1/\delta". For all "n\in\mathbb{N}", if "n>N" then "1/n<1/N<\delta" and hence,

"|\frac{f(c+1/n)-f(c)}{1/n}-k|<\varepsilon", that is, "|n({f(c+\frac{1}{n})-f(c)})-k|<\varepsilon".

This means that "\lim\limits_{n\to +\infty}n({f(c+\frac{1}{n})-f(c)})=k".


d) Let "f(x)=\cos\frac{2\pi}{x}", if "x\ne 0", else "f(x)=1". Consider "c=0".

Then

"\lim\limits_{h\to 0}\frac{f(c+h)-f(c-h)}{2h}=\lim\limits_{h\to 0}\frac{\cos\frac{2\pi}{h}-\cos\frac{2\pi}{(-h)}}{2h}=0" and

"\lim\limits_{n\to +\infty}n({f(c+\frac{1}{n})-f(c)})=\lim\limits_{n\to +\infty}n({\cos 2\pi n-1})=0"

We see that both conditions b) and c) are met, but "f(x)" is not differentiable at "x=0". It is not even continuous there, since the sequence "f(c+\frac{1}{2n})=\cos\pi n=(-1)^n" has no limit.